OpenStudy (anonymous):

what is the aproximate volume of a cone shaped party hat with a base whose circumference is 25.12 inches and height is 8 inches

OpenStudy (anonymous):

volume of cone is 1/3 pi r^2 h 2pi r=25.12 r=25.12/2pi r=4.0

OpenStudy (anonymous):

vol= 1/3 * pi* (4.0)^2 *8 =134.1cubic inches

OpenStudy (anonymous):

oh bt i have multiple coice and a.401.92 b.33.49 c.133.97 d.535.89 so whivh one would it be

OpenStudy (anonymous):

C

OpenStudy (anonymous):

k thanx

OpenStudy (anonymous):

so beens its c right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how is it c though if u told me that its 134.1

OpenStudy (anonymous):

if you HAVE to choose...there will be a margin of error when calculating

OpenStudy (anonymous):

whenever u round of things and continue calculatinbg with the figures that happens...pi by it sely is a rounded figure...i used 3.142

OpenStudy (anonymous):

oh okay now i get well thanx .

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you try to answer my friends question by smiley75 because we have the same questions